# DualStreamFoveaNet: A Dual Stream Fusion Architecture with Anatomical Awareness for Robust Fovea Localization

### v4

@@ 6 Conclusions
Accurate detection of the fovea is crucial for diagnosing retinal diseases. While anatomical structures outside the fovea, such as blood vessel distribution, are related to the fovea, few deep learning approaches exploit them to enhance fovea localization performance. In this paper, we propose a novel architecture, DualStreamFoveaNet (DSFN), which effectively fuses features from the retina and corresponding vessel distribution to achieve robust fovea localization. The DSFN comprises a two-stream encoder for multi-cue fusion and a decoder for generating fovea localization maps. Specifically, We introduce the Bilateral Token Incorporation (BTI) with learnable tokens to enhance the efficiency of transformer-based fusion from both fundus and vessel images. Comprehensive experiments demonstrate the advantages of using DSFN, including more accurate localization results, insensitivity to diseased images, and low computational cost. Our proposed architecture sets a new state-of-the-art on two public datasets (Messidor and PALM) and one large-scale private dataset (Tisu) with metrics ranging from 1/4 $R$ to $2R$. Additionally, it outperforms other methods in cross-dataset experiments, showcasing superior generalization capacity.

### v5

@@ 6 Conclusions
Summary. Accurate detection of the fovea is crucial for diagnosing retinal diseases. While anatomical structures outside the fovea, such as blood vessel distribution, are related to the fovea, few deep learning approaches exploit them to enhance fovea localization performance. In this paper, we propose a novel architecture, DualStreamFoveaNet (DSFN), which effectively fuses features from the retina and corresponding vessel distribution to achieve robust fovea localization. The DSFN comprises a two-stream encoder for multi-cue fusion and a decoder for generating fovea localization maps. Specifically, We introduce the Bilateral Token Incorporation (BTI) with learnable tokens to enhance the efficiency of transformer-based fusion from both fundus and vessel images. Comprehensive experiments demonstrate the advantages of using DSFN, including more accurate localization results through enhanced Multi-Cue Fusion, insensitivity to diseased images, and computational efficiency. Furthermore, our proposed architecture sets a new state-of-the-art on two public datasets (Messidor and PALM) and one large-scale private dataset (Tisu) with metrics ranging from 1/4 $R$ to $2R$. Additionally, it outperforms other methods in cross-dataset experiments, showcasing superior generalization capacity and robustness across datasets.
Limitations and Future Works. Although the promising performance of the proposed DSFN shown in our extensive experiments, the quality of the pseudo-mask of vessel structure still slightly affects the final fovea localization performance, as shown in Table 5. Compared to a fuzzy fovea region with an indistinguishable appearance from nearby textures, vessels are very distinct structures with high contrast. In this paper, the significant performance increase of DSFN relies on the vessels’ pseudo-masks generated by the TransUNet, which was trained on a small-scale vessel dataset. Recent studies [59, 60, 61] have specifically focused on vessel structure prediction, potentially providing more accurate pseudo-masks of vessels for our architecture in future research. Moreover, recently emerged Segment Anything Model (SAM)-based segmentation foundation models [62, 63], and their medical fine-tuned versions [64, 65], have shown significant disadvantages in predicting regions with no distinct boundaries and textures. Therefore, our design (i.e., utilizing global vessel structures to explicitly assist fovea localization) remains promising in this area.
Furthermore, the design of DSFN can be utilized as a general architecture for global feature integration and anatomical structure-guided segmentation tasks, as the input of the main and second streams can be any related images. In future work, we may upgrade DSFN for more medical imaging areas. In clinical practice, there are other tasks where even experienced clinicians require anatomical structure-based disease diagnosis. For example, in lung cancer screening, CT scans are commonly used for early detection. With the help of rib structures, doctors can better locate and observe lung lesions. In pelvic MRI, the structure of the pelvis can assist in the detailed evaluation of nearby pelvic organs and aid in diagnosing endometriosis and pelvic tumors.
Appendix
6.1 Notation Summary
All notions and abbreviations of this paper have been summarized in Table 7.
6.2 Augmentation Settings
For all input images, to reduce overfitting, we adopt a data augmentation strategy utilizing a PyTorch package, albumentations [66]. The details are shown as follows:
RandomResizedCrop (height = 512, width = 512, scale = (0.95, 1.05), ratio = (0.95, 1.05), p = 0.25), HorizontalFlip (p = 0.25), VerticalFlip (p = 0.25), ShiftScaleRotate (shift_limit = 0.0625, scale_limit = 0.05, rotate_limit = 45, p = 0.25), OneOf ([Blur (blur_limit = 5), GaussianBlur (blur_limit = 5), MedianBlur (blur_limit = 5), MotionBlur (blur_limit = 5)], p = 0.25), RandomBrightnessContrast (brightness_limit = 0.1, contrast_limit = 0.1, p = 0.25), OneOf ([CLAHE (clip_limit = 2), IAASharpen(), JpegCompression(), GaussNoise()], p = 0.25), CoarseDropout (p = 0.1).
Table 7: Notation Summary
Abbreviation Explanation
Bi-ViT Bilateral-ViT: an existing network [15] for fovea localization with bilateral feature fusion design
Bi-ViT/Lit The light version of Bilateral-ViT
BTI Bilateral Token Incorporation: The modules designed in this paper for globally bilateral feature fusion with high efficacy
CNN Convolutional Neural Networks
DSFN DualStreamFoveaNet: The two-stream architecture proposed in this paper
(G)FLOPs (Giga) Floating Point Operations
MF Multi-cue Features
MHSA Multi-Head Self-Attention
OD Optic Disc
ROI Region of Interest
RSU ReSidual U-blocks: The U-shape-based blocks utilized in the decoder of DSFN to integrate received features
TL and TF TokenLearner and TokenFuser
ViT Vision Transformer
Notation Description
$\text{F}^{\text{main}}_{\text{in}},\text{F}^{\text{satellite}}_{\text{in}}$ Input tensors from the main and satellite streams, respectively
$\text{F}_{\text{in}}$ Concatenated input feature
$\text{F}_{\text{attn}}$ Spatial attention map
$\text{F}^{\prime}_{\text{in}}$ Remapped input feature to generate $\text{T}_{\text{L}}$
$\text{T}_{\text{L}}$ Learned tokens
$\text{T}^{\prime}_{\text{L}}$ Intermediate token representation after MHSA layers
$\text{F}^{\prime\prime}_{\text{in}}$ Remapped input feature to generate $\text{F}_{\text{out}}$
$\text{F}_{\text{out}}$ Output features generated by $\text{T}^{\prime}_{\text{L}}$ and $\text{F}^{\prime\prime}_{\text{in}}$
$\text{F}^{\prime\text{main}}_{\text{out}},\text{F}^{\prime\text{satellite}}_{\text{out}}$ Output features equally split by $\text{F}_{\text{out}}$
$\text{F}^{\text{main}}_{\text{out}},\text{F}^{\text{satellite}}_{\text{out}}$ Final output features for the main and satellite streams, respectively